# Harmonic Loss Trains Interpretable AI Models

### v1

@@ Abstract
In this paper, we introduce harmonic loss as an alternative to the standard cross-entropy loss for training neural networks and large language models (LLMs). Harmonic loss enables improved interpretability and faster convergence, owing to its scale invariance and finite convergence point by design, which can be interpreted as a class center. We first validate the performance of harmonic models across algorithmic, vision, and language datasets. Through extensive experiments, we demonstrate that models trained with harmonic loss outperform standard models by: (a) enhancing interpretability, (b) requiring less data for generalization, and (c) reducing grokking. Moreover, we compare a GPT-2 model trained with harmonic loss to the standard GPT-2, illustrating that the harmonic model develops more interpretable representations. Looking forward, we believe harmonic loss has the potential to become a valuable tool in domains with limited data availability or in high-stakes applications where interpretability and reliability are paramount, paving the way for more robust and efficient neural network models.
Machine Learning, ICML
@@ 1 Introduction
In recent years, machine learning models have gained significant popularity, profoundly impacting various aspects of daily life. Consequently, it has become increasingly important to thoroughly understand the behavior of neural networks. One particularly intriguing characteristic of neural networks is their ability to generalize – empirical evidence shows that neural networks can perform well on unseen data not explicitly encountered during training (Novak et al., 2018). This remarkable ability stems from the networks’ capacity to learn generalizable representations and algorithms through training.
However, current models face three key challenges when it comes to generalization:
@@
(1) Lack of interpretability: Neural networks often lack interpretability. This opacity limits our understanding of how these models arrive at their conclusions, which is a critical issue in high-stakes applications like healthcare, finance, and autonomous systems. While mutiple research efforts have advanced our insight into the inner workings of LLMs (Bereska & Gavves, 2024), we are still far from fully explaining their outputs. Ultimately, we believe it is crucial to design systems that are interpretable by design. Without interpretability, it becomes challenging to diagnose errors, ensure fairness, or build trust in a model’s decisions.
@@
(2) Low data efficiency: Generalization often requires vast and diverse training data. This raises a critical question: can models generalize effectively with less data? This issue is especially relevant in domains where data availability is scarce, such as rare disease diagnosis, low-resource language processing, or specialized scientific fields like materials science or drug discovery. Previous approaches for improving neural network generalization include efficient data sampling (Li et al., 2024a) and modifications to the training procedure to accelerate training (Wang et al., 2024). However, these methods focus on optimizing existing training procedures rather than addressing the core issues in model design.
@@
(3) Delayed generalization (grokking): Models sometimes experience a phenomenon known as “grokking,” (Power et al., 2022; Liu et al., 2021) where there is a noticeable delay between the convergence of the training loss and the convergence of the test loss. This gap is problematic because: (i) it complicates determining the optimal point to stop training in order to achieve generalization, and (ii) it necessitates extended computation time and resources to continue training until grokking occurs.
@@
As the saying goes, “The devil is in the details.” We attribute these three challenges in part to the widespread use of cross-entropy loss (for classification) and propose harmonic loss as an alternative. Harmonic loss has two desirable mathematical properties that enable faster convergence and improved interpretability: (1) scale invariance, and (2) a finite convergence point, which can be interpreted as a class center. Through a comprehensive set of experiments, we demonstrate that models trained with harmonic loss outperform standard models in terms of reducing grokking, requiring less data for generalization, and enhancing interpretability. Furthermore, we compare a GPT-2 model trained with harmonic loss to the standard GPT-2 and show that the harmonic model develops more interpretable representations.
@@
The remainder of this paper is organized as follows: In Section 2, we review the relevant literature. Section 3 introduces the principles underlying harmonic loss and explains why it is preferable to cross-entropy loss in terms of generalization capabilities. Section 4 details a comprehensive set of experiments on algorithmic datasets, demonstrating that models trained with harmonic loss consistently outperform standard models. In Section 5, we demonstrate the performance of harmonic models on the vision task of MNIST digit classification. In Section 6, we extend our analysis to large models, illustrating that the advantages of harmonic loss also hold at scale. Finally, we conclude the paper in Section 7.
2 Related Works
Representations: In this paper, we aim to improve the interpretability of neural network representations. Numerous studies have shown that LLMs can form conceptual representations across spatial (Gurnee & Tegmark, 2023), temporal (Li et al., 2021), and color domains (Abdou et al., 2021). The structure of such representations includes one-dimensional concepts (Gurnee & Tegmark, 2023; Marks & Tegmark, 2023; Heinzerling & Inui, 2024; Park et al., 2024b), as well as multi-dimensional representations such as lattices (Michaud et al., 2024; Li et al., 2024c) and circles (Liu et al., 2022a; Engels et al., 2024). Recent works have also studied the representations developed during inference-time (Park et al., 2024a). While the structure of these representations often correlates with certain geometric patterns, significant unexplained variance frequently remains, complicating interpretability.
Mechanistic Interpretability: While increasing the number of parameters and the amount of data samples used to train neural networks has enhanced their capabilities, it has also made mechanistically interpreting these models more challenging. This line of work has been explored from two major directions: the circuit level (Michaud et al., 2024; Olah et al., 2020a; Olsson et al., 2022; Templeton et al., 2024), which aims to identify a submodule within an LLM responsible for a specific ability, and the representation level (Liu et al., 2022b; Ding et al., 2024; Zhong et al., 2024). Some works have made progress by designing interpretable systems through the decomposition of models into smaller modules (Olah et al., 2020a; Liu et al., 2023).
Loss Functions: Previous research has shown that loss functions can influence how a model learns to represent data, affecting its abilities in unique ways (Li et al., 2024b; Bosco et al., 2024). Novel loss functions have improved performance on specific tasks, though they often reduce ability in different settings (Bommidi et al., 2023; Seber, 2024). For instance, focal loss, dice loss, and Tversky loss have proven effective for image segmentation (Sudre et al., 2017; Demir et al., 2023; Salehi et al., 2017), but only focal loss is also effective for object detection (Lin, 2017). Luo et al. (2021) found that smoothly approximating non-smooth Hinge loss in Support Vector Machines (SVMs) improved the convergence rate of optimization.
@@ 3 Harmonic Loss
Figure 1: Cross-entropy loss versus harmonic loss (ours). (a) Definitions. Cross-entropy loss leverages the inner product as the similarity metric, whereas the harmonic loss uses Euclidean distance. (b) Toy case 1 with two points (classes). Both the harmonic loss and the $l_{2}$ weight norm converge faster for the harmonic loss. (c) Toy case 2 with five points (classes). Harmonic loss can pick out the red point in the middle. By contrast, the cross-entropy loss cannot, since the red point is not linearly separable from other points. The weight matrices are also more interpretable with harmonic loss than with cross-entropy loss.
@@
Cross-entropy loss: Logits $\bm{y}$ are defined as the matrix-vector multiplication, i.e., $\bm{y}=\bm{W}^{T}\bm{x}\in\mathbb{R}^{V}$ (ignoring biases), or
$y_{i}=\bm{w}_{i}\cdot\bm{x},$ (1)
where $\bm{w}_{i}$ is the $i^{\rm th}$ column of $\bm{W}$. Probability $\bm{p}$ can be obtained by applying softmax to $\bm{y}$, i.e.,
@@
Harmonic loss: The harmonic logit $\bm{d}$ is the $l_{2}$ distance between $\bm{w}_{i}$ and $\bm{x}$, i.e.,
$d_{i}=||\bm{w}_{i}-\bm{x}||_{2}.$ (3)
We interpret $\bm{w}_{i}$ as keys and $\bm{x}$ as a query, so smaller $d_{i}$ means a higher probability of $p_{i}$. We define harmonic max (harmax) as
@@
A reasonable choice of $n$ is $n\sim\sqrt{D}$, where $D$ represents the intrinsic dimensionality of the underlying data. In LLMs, $D$ could be approximated as $D\approx d_{\rm embed}$, where $d_{\rm embed}$ is the embedding dimension. This approximation arises from considering an embedding initialized from a $D$-dimensional Gaussian distribution. The squared distance between two points, normalized by the number of dimensions $D$, is on the order of $1\pm O(1/\sqrt{D})$. To ensure that the harmonic distance $\left[1\pm O(1/\sqrt{D})\right]^{n}$ remains constant as we scale $D$, we require $n\sim\sqrt{D}$, since $\lim_{x\rightarrow\infty}(1+x^{-1})^{x}=e$.
@@
Benefits of harmonic loss: From these two toy cases, we understand the advantages of harmonic loss: (1) nonlinear separability: in case 2, the red dot can be classified correctly even though it is not linearly separable. (2) fast convergence: The fact that the converging point is finite leads both to faster loss decay, and plateauing (non-diverging) $l_{2}$. (3) scale invariance: Harmonic loss is scale-invariant, i.e., $d_{i}\to\alpha d_{i}$ leaves $p_{i}$ (hence loss) invariant, whereas $y_{i}\to\alpha y_{i}$ would produce a different cross-entropy loss. (4) interpretability: the weight vectors correspond to class centers.
The rest of the paper explores the use of harmonic loss in various applications: algorithmic tasks in Section 4, MNIST in Section 5, and language modeling in Section 6.
4 Algorithmic Datasets
Algorithmic tasks are good benchmarks for network interpretability since they are well-defined mathematically. However, training neural networks on these tasks is non-trivial due to the observation of grokking (delayed generalization) (Power et al., 2022) and the existence of multiple algorithms (Zhong et al., 2024), etc. We will show that harmonic models can learn better representations, are more data-efficient, and experience less grokking.
@@ 4.1 Models and Datasets
    Modular Addition: Given two integers $x$ and $y$, the model is trained to predict $(x+y)\;\textrm{mod}\;31$.
@@ 4.2 Representation Faithfulness
Figure 2 shows the plot of the top two principal components of the models’ embeddings for MLP tasks. A complete visualization is available in Figure 2. We show the embedding visualization of transformers in Appendix A. Overall, harmonic loss representations are cleaner and more organized than their cross-entropy counterparts. We found near-perfect circle representations for the modular addition task, a clear tower-like structure for tree learning, and neat clusters for permutation composition.
We examine the representations task by task:
(1) In-context learning: We observe that the representations obtained from standard models are either imperfect lattices or exhibit unexplained variance in higher dimensions, whereas harmonic models almost always perfectly ($100\%$) recover the underlying 2D lattice structure regardless of the random seed.
(2) Modular addition: Harmonic MLPs consistently recover a purely 2D circular representation in almost all runs, whereas the standard MLP often fails to identify the circular structure. While the harmonic transformer has a similar success rate to the standard transformer in constructing circles, the explained variance captured by the first two principal components is generally much higher, indicating that harmonic models discover more compact representations with fewer uninterpretable components.
(3) Equivalence classes: Both standard and harmonic models are able to identify the underlying groups with high variance. However, we note that standard models’ representation tends to be more “elongated”, or not completely grouped, compared to its harmonic counterpart. This could be attributed to the fact that cross-entropy loss does not have an incentive to reduce irrelevant variations to zero.
(4) Genealogy learning: Harmonic MLP is the only model that successfully recovers the underlying tree representation.
(5) Permutation composition: The harmonic MLP generally produces better-separated clusters. A particularly clean representation that appears multiple times contains 6 clusters of 4 permutations, where each cluster is a coset of the subgroup $\langle e,(12)(34),(13)(24),(14)(23)\rangle$ or one of its conjugates. In the harmonic transformer, permutations commonly organize into 4 clusters that are cosets of $\langle e,(13),(14),(34),(134),(143)\rangle$ or one of its conjugates, subgroups isomorphic to $S_{3}$ (one element, in this case $2$, never permutes).
@@ 4.3 Data Efficiency in Training
Figure 3: (a) Cumulative explained variance as a function of principal components (median over 20 seeds). Harmonic representations are more compact than standard counterparts. (b) Test Accuracy as a function of Training Fraction. Harmonic models generalize faster with less data than standard counterparts. (c) Epochs to Test Accuracy $>$ 0.9 vs Epochs to Train Accuracy $>$ 0.9 for 20 consecutive times. $y=x$ line represents no grokking, where train and test accuracy improve simultaneously. Points closer to the y-axis indicate a greater degree of grokking. Results from 20 different random seeds are plotted, and the runs that were not able to achieve 90% accuracy were omitted.
Humans can learn from surprisingly few samples, but neural networks usually require much more data to learn. As a result, we want models that can learn efficiently from data. Figure 3(b) shows the test accuracy as a function of train data fraction for our synthetic experiments, indicating how much data is necessary in order for the model to be generalizable. We observe that harmonic models require comparable or much less amount of data to generalize, compared to their cross-entropy counterparts. Such improvement is especially notable for in-context learning, where harmonic models generalize nearly immediately.
@@ 4.4 Reduced Grokking
Grokking refers to the phenomenon of delayed generalization (Power et al., 2022): for example, it takes $10^{3}$ steps to reach perfect accuracy on the training data, but it takes $10^{5}$ steps to generalize to the test data. Grokking is a pathological phenomenon that we want to avoid (Liu et al., 2022b). We find that harmonic loss overall reduces grokking, as seen in Figure 3(c). Points on the $y=x$ line represent models which trained without grokking, with train and test accuracy improving together. This improvement is particularly evident in learning modular addition and permutation composition: while the standard MLP exhibits severe grokking, most data points for the harmonic MLP lie much closer to the $y=x$ line.
@@ 4.5 Case Study: Modular Addition
Figure 4: Case study on modular addition. Standard MLP trained for modular addition without weight decay often fails to generalize. Generalization is only achieved with the addition of strong weight decay; however, (a) significant grokking occurs, and (b) while the first two principal components form an approximate circle, they explain far less than the total variance. In contrast, the harmonic model trained for modular addition generalizes quickly without grokking. Moreover, the embedding forms a perfect 2D circle. EV in the plot represents the explained variance by the first two principal components of the embedding.
@@
The better formation of a circle and improved generalization in harmonic MLP can be attributed to the properties of harmonic loss, as explained in Section 3. To drive the probability to 1, the standard cross-entropy loss requires driving the representation to infinity – i.e., making the logit infinite. In contrast, harmonic loss achieves this by driving the harmonic logit to zero, which is easily accomplished by learning $\bm{w}_{i}=\bm{x}$ in Equation 3. The existence of such a finite converging point results in (a) faster convergence, (b) better generalization, and (c) more interpretable representations.

### v2

@@ Abstract
In this paper, we introduce harmonic loss as an alternative supervisory signal for training neural networks and large language models (LLMs). Harmonic loss differs from standard cross-entropy loss by (a) replacing the usual SoftMax normalization with a scale-invariant HarMax function and (b) computing logits via Euclidean distance rather than a dot product. Harmonic loss enables improved interpretability and faster convergence, owing to its scale invariance and finite convergence point by design, which can be interpreted as a class center. We first validate the performance of harmonic models across algorithmic, vision, and language datasets. Through extensive experiments, we demonstrate that models trained with harmonic loss perform better than standard models by: (a) enhancing interpretability, (b) requiring less data for generalization, and (c) reducing grokking. Moreover, we compare a GPT-2 model trained with harmonic loss to the standard GPT-2, illustrating that the harmonic model develops more interpretable representations. Looking forward, we believe harmonic loss may become a valuable tool in domains with limited data availability or in high-stakes applications where interpretability and reliability are paramount, paving the way for more robust and efficient neural network models.
@@ 1 Introduction
As machine learning models become powerful, it has become increasingly important to thoroughly understand the behavior of neural networks. One particularly intriguing characteristic of neural networks is their ability to generalize—empirical evidence shows that neural networks can perform well on unseen data not explicitly encountered during training [1]. This remarkable ability stems from the networks’ capacity to learn generalizable representations and algorithms through training. However, current models face three key challenges when it comes to generalization:
@@
(1) Lack of interpretability: Neural networks often lack interpretability, which is a critical issue in high-stakes applications like healthcare, finance, and autonomous systems. While multiple research efforts have advanced our insight into inner workings of LLMs [2], we are still far from fully explaining their outputs. Ultimately, it is crucial to design systems that are interpretable by design. Otherwise, it is challenging to diagnose errors, ensure fairness, or build trust in a model’s decisions.
@@
(2) Low data efficiency: Generalization often requires vast and diverse training data. This raises a critical question: can models generalize effectively with less data? This issue is especially relevant in domains where data availability is scarce, such as rare disease diagnosis or specialized scientific fields. Previous approaches for improving neural network generalization include efficient data sampling [3] and modifications to the training procedure to accelerate training [4]. However, these methods focus on optimizing existing training procedures rather than addressing the core issues in model design.
@@
(3) Delayed generalization (grokking): Models sometimes experience a phenomenon known as “grokking,” [5, 6] where there is a noticeable delay between the convergence of the training loss and the convergence of the test loss. This gap is problematic because: (i) it complicates determining the optimal point to stop training in order to achieve generalization, and (ii) it necessitates extended computation time and resources to continue training until grokking occurs.
@@
As the saying goes, “The devil is in the SoftMax.” We attribute these three challenges in part to the widespread use of the SoftMax function in cross-entropy loss (for classification) and propose harmonic loss as an alternative. Harmonic loss has two desirable mathematical properties that enable faster convergence and improved interpretability: (1) scale invariance, and (2) a finite convergence point, which can be interpreted as a class center. Through comprehensive experiments, we show that models trained with harmonic loss reduce grokking, require less data for generalization, and enhance interpretability compared to standard models. Furthermore, we compare a GPT-2 model trained with harmonic loss to the standard GPT-2 and show that the harmonic model develops more interpretable representations.
@@
The remainder of this paper is organized as follows: Section 2 introduces the principles underlying harmonic loss and explains why it is preferable to cross-entropy loss in terms of generalization and interpretability. Section 3 details a comprehensive set of experiments on algorithmic datasets, illustrating that models trained with harmonic loss have numerous desirable properties that are absent in standard models. In Section 4, we demonstrate the performance of harmonic models on the vision task of MNIST digit classification. In Section 5, we extend our analysis to large models, illustrating that the advantages of harmonic loss also hold at scale. We present ablation experiments in Section 6. We review the relevant literature in Section 7, and conclude the paper in Section 8.
@@ 2 Harmonic Loss
Figure 1: Cross-entropy loss versus harmonic loss (ours). (a) Definitions. Cross-entropy loss leverages the inner product as the similarity metric, whereas the harmonic loss uses Euclidean distance. (b) Toy case 1 with two points (classes). Both the loss and $l_{2}$ weight norm converge faster for the harmonic loss. (c) Toy case 2 with five points (classes). Harmonic loss can pick out the red point in the middle. By contrast, the cross-entropy loss cannot, since the red point is not linearly separable from other points. Weight matrices are also more interpretable with harmonic loss than with cross-entropy loss.
@@
Cross-entropy loss: Logits $\bm{y}$ are defined as the matrix-vector multiplication, i.e., $\bm{y}=\bm{W}^{T}\bm{x}\in\mathbb{R}^{V}$ (ignoring biases), or $y_{i}=\bm{w}_{i}\cdot\bm{x}$, where $\bm{w}_{i}$ is the $i^{\rm th}$ column of $\bm{W}$. Probability $\bm{p}$ can be obtained by applying SoftMax to $\bm{y}$, i.e.,
@@
Harmonic loss: The harmonic logit $\bm{d}$ is the $l_{2}$ distance between $\bm{w}_{i}$ and $\bm{x}$, i.e., $d_{i}=||\bm{w}_{i}-\bm{x}||_{2}.$ We interpret $\bm{w}_{i}$ as keys and $\bm{x}$ as a query, so smaller $d_{i}$ means a higher probability of $p_{i}$. We define harmonic max (HarMax) as
@@
A reasonable choice of $n$ is $n\sim\sqrt{D}$, where $D$ represents the intrinsic dimensionality of the underlying data. In LLMs, $D$ could be approximated as $D\approx d_{\rm embed}$, where $d_{\rm embed}$ is the embedding dimension. This approximation arises from considering an embedding initialized from a $D$-dimensional Gaussian distribution. The squared distance between two points, normalized by the number of dimensions $D$, is on the order of $1\pm O(1/\sqrt{D})$. To ensure that the harmonic distance $\left[1\pm O(1/\sqrt{D})\right]^{n}$ remains constant as we scale $D$, we require $n\sim\sqrt{D}$, since $\lim_{x\rightarrow\infty}(1+x^{-1})^{x}=e$. We also show the empirical impact of the exponent on the learned representations in Appendix E.
@@
Benefits of harmonic loss: From these two toy cases, we understand the advantages of harmonic loss: (1) nonlinear separability: in case 2, the red dot can be classified correctly even though it is not linearly separable. (2) fast convergence: The fact that the converging point is finite leads both to faster loss decay, and plateauing (non-diverging) $l_{2}$. (3) scale invariance: Harmonic loss is scale-invariant, i.e., $d_{i}\to\alpha d_{i}$ leaves $p_{i}$ (hence loss) invariant, whereas $y_{i}\to\alpha y_{i}$ would produce a different cross-entropy loss. (4) interpretability: the weight vectors correspond to class centers. We present the formal proof of these properties in Appendix G.
Notes on interpretability: Measuring interpretability is inherently challenging in the absence of ground-truth representations. Hence, we propose two principled indicators of interpretability throughout the paper: (1) Compression: Sparse, low-dimensional representations enhance interpretability by concentrating semantics. We measure this via cumulative explained variance in PCA projections. (2) Geometry: In general models, we hypothesize that parallelogram-like units with multiple one-dimensional semantic directions enable compositional reasoning; This enables vector arithmetic such as man – woman = king – queen, and supports faithful feature attribution. We measure this via parallelogram loss in Section 5.
3 Algorithmic Experiments
Algorithmic tasks are good benchmarks for interpretability since they are well-defined mathematically. However, training neural networks on these tasks is non-trivial due to grokking (delayed generalization) [5] and the existence of multiple algorithms [7], etc. We will show that harmonic models learn better representations, are more data-efficient, and exhibit less grokking.
@@ 3.1 Models and Datasets
    Modular Addition: Given two integers $x,y$, the model is trained to predict $(x+y)\;\textrm{mod}\;31$.
@@ 3.2 Representation Faithfulness
Figure 3: (a) Cumulative explained variance as a function of principal components (mean over 20 seeds). Harmonic representations are more compact than standard counterparts. (b) Test Accuracy as a function of Train Fraction (mean over 3 seeds). Harmonic models generalize faster with less data than standard counterparts. (c) Epochs to Test Accuracy $>$ 0.9 vs Epochs to Train Accuracy $>$ 0.9 for 20 consecutive epochs. $y=x$ line represents no grokking, and points closer to the y-axis indicate more grokking. Results from 20 different random seeds are plotted, and the runs that were not able to achieve 90% accuracy were omitted. We present the plots for all tasks in Appendix F.
Figure 2 shows the plot of the top two principal components of the models’ embeddings for MLP tasks. We show the complete embedding visualization for all tasks in Appendix A. Overall, harmonic loss representations are cleaner and more organized than their cross-entropy counterparts. We found near-perfect circle representations for the modular addition task, a clear tower-like structure for tree learning, and neat clusters for permutation composition. We examine the representations task by task:
1. In-context Learning: Standard models’ representations are either imperfect lattices or exhibit unexplained variance in higher dimensions, whereas harmonic models almost always perfectly ($100\%$) recover the underlying 2D lattice structure across different random seeds.
2. Modular Addition: Harmonic MLPs consistently recover a perfect 2D circular representation in almost all runs, whereas tstandard MLPs often fail to do so. Harmonic transformer has a similar success rate to the standard transformer in constructing circles, but the explained variance captured by the first two principal components is generally much higher, indicating that harmonic models discover more compact representations with fewer uninterpretable components.
3. Equivalence Classes: While both standard and harmonic models are able to identify the underlying groups, standard models’ representation tends to be more “elongated", or not completely grouped, compared to its harmonic counterpart. This could be attributed to the fact that cross-entropy loss does not have an incentive to reduce irrelevant variations to zero.
4. Genealogy Learning: Only Harmonic MLP recovers the underlying tree representation.
5. Permutation Composition: Harmonic MLP generally produces better-separated clusters. A particularly clean representation that appears multiple times contains 6 clusters of 4 permutations, where each cluster is a coset of the subgroup $\langle e,(12)(34),(13)(24),(14)(23)\rangle$ or one of its conjugates. In the harmonic transformer, permutations commonly organize into 4 clusters that are cosets of $\langle e,(13),(14),(34),(134),(143)\rangle$ or one of its conjugates, subgroups isomorphic to $S_{3}$ (one element, in this case $2$, never permutes).
@@ 3.3 Data Efficiency in Training
Figure 3(b) shows the test accuracy as a function of train data fraction for our synthetic experiments, indicating how much data is necessary in order for the model to be generalizable. We observe that harmonic models require comparable or much less amount of data to generalize, compared to their cross-entropy counterparts. Such improvement is especially notable for in-context learning, where harmonic models generalize nearly immediately.
@@ 3.4 Reduced Grokking
Grokking refers to the phenomenon of delayed generalization [5]: for example, it takes $10^{3}$ steps to reach perfect accuracy on the training data, but it takes $10^{5}$ steps to generalize to the test data. Grokking is a pathological phenomenon that we want to avoid [8]. We find that harmonic loss overall reduces grokking, as seen in Figure 3(c). Points on the $y=x$ line represent models which trained without grokking, with train and test accuracy improving together. This improvement is particularly evident in learning modular addition and permutation composition: while the standard MLP exhibits severe grokking, most data points for the harmonic MLP lie much closer to the $y=x$ line.
@@ 3.5 Case Study: Modular Addition
Figure 4: Left: Case study on modular addition. Standard MLP trained for modular addition without weight decay often fails to generalize. Generalization is only achieved with the addition of strong weight decay; however, (a) significant grokking occurs, and (b) while the first two principal components form an approximate circle, they explain far less than the total variance. In contrast, the harmonic model trained for modular addition generalizes quickly without grokking. Moreover, the embedding forms a perfect 2D circle. EV in the plot represents the explained variance by the first two principal components of the embedding. Right: Visualization of model weights trained for MNIST. Yellow cells show values less than $0.01$. Both models achieved $\approx 92.5\%$ test accuracy.
@@
The better formation of a circle and improved generalization in harmonic MLP can be attributed to the properties of harmonic loss, as explained in Section 2. To drive the probability to 1, the standard cross-entropy loss requires driving the representation to infinity—i.e., making the logit infinite. In contrast, harmonic loss achieves this by driving the harmonic logit to zero, which is easily accomplished by learning $\bm{w}_{i}=\bm{x}$ in Equation 2. The existence of such a finite converging point results in (a) faster convergence, (b) better generalization, and (c) more interpretable representations.